OpenStudy (anonymous):

A comet that passes through the solar system just once has a path that is modeled by a conic section. Astronomers discover a comet whose path is modeled by the following equation, with the Sun at one focus. What type of conic section is the comet's path? (x^2)/(2500) - (y^2)/(37500) =1

OpenStudy (turingtest):

circle:\[x^2+y^2=r^2\]ellipse:\[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\]hyperbola:\[\frac{x^2}{a^2}-\frac{y^2}{b^2}=1\]pick one!

OpenStudy (anonymous):

im not sure which one

OpenStudy (turingtest):

look very hard to find the differences between these three choices only one of those traits appears in your formula

OpenStudy (anonymous):

oh ok thanks lol i cant belive i missed that.